A 0.020 M solution of a weak acid HA has a pH of 3.50. What is the percent ionization of HA in the solution?

For the solution described above, what is the Ka?
Answer #1

use:
pH = -log [H+]
3.5 = -log [H+]
[H+] = 3.162*10^-4 M
HA dissociates as:

HA          ----->     H+   + A-
2*10^-2                 0         0
2*10^-2 - x               x         x

Here:
x = [H+] = 3.162*10^-4 M

1)
% dissociation = x * 100 / initial concentration
= (3.162*10^-4) * 100 / 0.020
= 1.58 %
Answer: 1.58 %

2)

Ka = [H+][A-]/[HA]
Ka = x*x/(c-x)
Ka = 3.162*10^-4*3.162*10^-4/(0.02-3.162*10^-4)
Ka = 5.08*10^-6
Answer: 5.08*10^-6